# Memory Decoder: A Pretrained, Plug-and-Play Memory for Large Language Models

### v1

@@ G.1 Extreme Sparsity and Concentration
G.2 Scale-Dependent Behavior
Model scale dramatically affects $k$-NN distribution quality. GPT-2-Small (117M) produces distributions marginally different from its LM outputs (top-1 probability  50%), while GPT-2-Large (1.5B) generates radically sparse distributions with 93.48% average top-1 probability—a 67% relative increase over its baseline.
@@ G.2 Scale-Dependent Behavior
Table 19: Perplexity with Memory Decoder (125M) trained using $k$-NN distributions from different source models
@@
G.3 Domain Adaptation Effects
@@
Appendix H Alternative Loss Functions for Imitating $k$-NN Distributions
H.1 Failed Approaches
@@ H.1.1 Focal Loss
H.1.2 Jensen-Shannon Divergence
@@
H.1.3 Bi-directional Logits Difference (BiLD)
@@
H.1.4 Explicit Sparsity Penalty

### v2

@@ C.1 Extreme Sparsity and Concentration
C.2 Scale-Dependent Behavior
Model scale dramatically affects $k$-NN distribution quality. GPT-2-small (124M) produces distributions marginally different from its LM outputs (top-1 probability 50%), while GPT-2-Large (1.5B) generates radically sparse distributions with 93.48% average top-1 probability—a 67% relative increase over its baseline.
@@ C.2 Scale-Dependent Behavior
Table 11: Perplexity with Memory Decoder (124M) trained using $k$-NN distributions from different source models
@@
C.3 Domain Adaptation Effects
@@
Appendix D Alternative Loss Functions for Imitating $k$-NN Distributions
D.1 Failed Approaches
@@ D.1.1 Focal Loss
D.1.2 Jensen-Shannon Divergence
@@
D.1.3 Bi-directional Logits Difference (BiLD)
@@
D.1.4 Explicit Sparsity Penalty